# How does Architecture Influence the Base Capabilities of Pre-trained Language Models? A Case Study Based on FFN-Wider Transformer Models

### v1

@@ Abstract
Pre-trained language models have been proven to possess strong base capabilities, which not only excel in in-distribution language modeling but also show powerful abilities in out-of-distribution language modeling, transfer learning and few-shot learning. Unlike existing work focusing on the influence of scale on base capabilities, our work examines the influence of architecture on those. Specifically, our concern is: How does architecture influence the base capabilities of pre-trained language models? In this work, we attempt to explain and reverse the decline in base capabilities caused by the architecture of FFN-Wider Transformers, seeking to provide some insights. Through analysis, we found the contribution ratio of Multi-Head Attention (a combination function) to pre-trained language modeling is a key factor affecting base capabilities. FFN-Wider Transformers reduce the contribution ratio of this combination function, leading to a decline in base capabilities. We confirmed this by experiments and proposed Combination Enhancement Architecture (CEA) to address the decline in base capabilities of such models. Significantly, we extended our explanation and CEA to Mixture of Experts (MoE) architecture Transformers, which also alleviated their decline in base capabilities to some extent, proving our work can offer useful guidance for architecture analysis, architecture improvement and architecture design.
How does Architecture Influence the Base Capabilities of
Pre-trained Language Models? A Case Study Based
on FFN-Wider Transformer Models
Xin Lu,   Yanyan Zhaothanks:     Email corresponding.,   Bing Qin Research Center for Social Computing and Information Retrieval Harbin Institute of Technology, China {xlu, yyzhao, qinb}@ir.hit.edu.cn
@@ 1 Introduction
Recent research has discovered that pre-trained language models possess strong base capabilities Radford et al. (2018); Devlin et al. (2019); Brown et al. (2020); OpenAI (2023). They can not only address in-distribution language modeling which is usually their pre-training objective, but also unexpectedly excel in out-of-distribution language modeling, transfer learning, few-shot learning, etc. This has attracted the attention of many researchers.
@@
Figure 1: The results of various architecture models at two different scales illustrate that: 1) Under similar pre-training performance, the FFN-Wider BERT models demonstrate a noticeable decline in base capabilities (such as out-of-distribution language modeling and transfer learning) compared to the vanilla BERT models. 2) The FFN-Wider BERT models with our Combination Enhanced Architecture (CEA) successfully reverse the decline in base capabilities.
Figure 2: Illustration showing that the synchronous improvement in model base capability as the actual contribution ratio of the Outer-FFN layer (a transformation function) decreases, that is, the actual contribution ratio of the MHA layer (a combination function) increases. This reveals a key factor affecting model’s base capabilities.
In this process, much attention has been focused on analyzing scale, leading to the formulation of compelling scaling laws Kaplan et al. (2020); Hoffmann et al. (2022) that drive the trend of enhancing base capability by increasing parameter numbers, data volume and training tokens. In contrast, the impact of architecture has not received sufficient attention. According to the basic principle of inductive bias in machine learning, model architecture is also a crucial factor affecting base capabilities, and its impact could be equally decisive.
Some studies have already noted the significant influence of architecture on the base capabilities. For instance,  (Tay et al., 2023) have observed considerable differences in base capabilities among various Transformer architecture variants. Some variants, though larger in scale and more powerful in pre-training performance than vanilla Transformers, exhibit significantly reduced performance in downstream tasks. This suggests different architecture models vary greatly in converting pre-training performance into base capabilities. Simply increasing scale does not resolve all issues, and exploring the impact of architecture on base capabilities is crucial.
@@
Specifically, we focus on the Transformer with wider FFN (Feed-Forward Network) layers, finding that this architecture change leads to a significant decrease in base capabilities compared to the vanilla Transformer. As shown in Figure 1, under similar pre-training performance, the FFN-Wider BERT models exhibit a noticeable decline in both out-of-distribution language modeling and downstream tasks fine-tuning compared to the vanilla BERT models. We believe such a simple change, leading to significant differences in base capabilities, makes for a good object of study in exploring how architecture impacts base capabilities.
@@
We firstly attempted to explain this change in base capabilities. Through analysis, we concluded that the MHA (Multi-Head Attention) layer in the Transformer is a combination function, and the FFN (Feed-Forward Network) layer is a transformation function, with the former being a focused expression of the combinability of language. During their contribution to pre-trained language modeling, the actual contribution ratio of the MHA layer, as a combination function, is a key factor affecting the model’s base capabilities. The FFN-Wider Transformer models directly enhance the FFN layer, indirectly reducing the combination function’s actual contribution ratio to pre-trained language modeling, thereby leading to a significant decline in base capabilities.
@@
To validate our explanation for it, we designed a Combination Adjustable Architecture (CAA), as depicted in Figure 2(a). This architecture bifurcates a wider FFN into two parts with adjustable width ratios: one remains in its original position as a transformation function, known as the Outer-FFN, and the other is relocated within the MHA layer, transformed through a special design into an Inner-FFN that solely enhances the combination function. We controlled the width ratio of the Outer-FFN, reducing it gradually from 100% to 0%. In Figure 2(b), we observed that the actual contribution ratio of the Outer-FFN progressively decreased, indicating a corresponding increase in the actual contribution ratio of the MHA layer. In Figure 2(c), we observed a gradual improvement in base capabilities. These reveal a key phenomenon that confirms our explanation: as the actual contribution ratio of the MHA layer (a combination function) increases, there is a general synchronous improvement in the model’s base capabilities.
@@
Subsequently, we identified the optimal width ratio of two parts of FFN for the Combination Adjustable Architecture (CAA) and established it as the Combination Enhanced Architecture (CEA), and comparing this architecture with the original FFN-Wider Transformer. As shown in Figure 1, the new architecture BERT models exhibit a comprehensive enhancement in base capabilities compared to the original FFN-Wider BERT models, nearly reaching the vanilla BERT model level, further validating our explanation. Ultimately, we conducted extensive analysis and experiments on various scales of BERT and GPT models, with all results robustly supporting our explanation.
Importantly, we also noticed that existing work has observed a similar decline in base capabilities in the Mixture of Experts (MoE) architecture Transformers. We applied our explanation and CEA to MoE Transformer models as well, achieving improvements in base capabilities, proving the applicability of our work to other architectures. Furthermore, this suggests that the actual contribution ratio of the MHA layer (a combination function) is likely a universal factor affecting model’s base capabilities, providing valuable insights for architecture analysis, improvement and design.
@@ 2.1 Base Capabilities
Transfer Learning   This is a recognized base capability of pre-trained language models. The works proposing GPT Radford et al. (2018) and BERT Devlin et al. (2019) have established the "pre-training and fine-tuning" transfer learning paradigm under the Transformer architecture.
Few-shot Learning   This is also a recognized base capability. Some works Radford et al. (2019); Brown et al. (2020) found large-scale pre-trained language models could complete many NLP tasks with no or only a few demonstrations, revealing their remarkable few-shot learning capabilities.
Additionally, recent large-scale pre-trained language models have demonstrated astonishing capabilities like following instructions Ouyang et al. (2022); OpenAI (2023). However, such capabilities often emerge only in super-large scale models. Since our work requires pre-training the models from scratch, we have not included these capabilities in our current work.
@@ 2.2 Evaluation Schemes
2.3 Experimental Settings
Unless specified otherwise, all experiments follow the unified settings described in this section.
The pre-training corpus is consistent with that of BERT Devlin et al. (2019), namely Wikipedia and BooksCorpus. We partition a portion of the corpus to serve as an in-distribution development set, and use it to align pre-training performance.
We conduct experiments on four specifications: BERT (H=128), BERT (H=768), GPT (H=128) and GPT (H=768). The specifications and pre-training procedures are detailed in Appendix B.
For out-of-distribution language modeling capability, we evaluate on the development set of Pile dataset Gao et al. (2020).
For transfer learning capability, we only evaluate BERT models on GLUE Wang et al. (2018), SuperGLUE Wang et al. (2019), HellaSwag Zellers et al. (2019), PIQA Bisk et al. (2020), OpenBookQA Mihaylov et al. (2018), ARC Easy & Challenge Clark et al. (2018) and WinoGrande Sakaguchi et al. (2021). The experimental settings are detailed in Appendix C.
For few-shot learning capability, we only evaluate GPT models. Limited by the maximum sequence length (128) of our pre-trained models, we conduct 0-shot and 1-shot experiments on HellaSwag, PIQA, OpenBookQA, ARC Easy & Challenge, WinoGrande, Winograd Levesque et al. (2012), COPA Roemmele et al. (2011) and StoryCloze Mostafazadeh et al. (2016). The experimental settings are detailed in Appendix D.
@@ 2.3 Experimental Settings
We found that, at the same level of pre-training performance, the Transformer models with wider FFN layers exhibit a noticeable decline in performance on most downstream tasks, indicating a deterioration in their base capabilities.
This presents us with a good research object. It evolved simply from the vanilla Transformer architecture, demonstrating a significant shift in base capabilities by altering a single clear variable. In the subsequent sections, we will conduct extensive research around the decline in the base capabilities of this architecture model.
Figure 3: Comparison of the base capabilities between FFN-Wider Transformers and Vanilla Transformers.
@@ 4.1 Combination and Transformation
Transformers consist of MHA and FFN layers. Considering a certain position in a sequence, the updated representation obtained after an MHA layer is a combination of the entire sequence context; whereas the updated representation obtained after an FFN layer is from a transformation of that position representation alone, context-insensitive. The diagrams are shown in Figure 4(a) and 4(b).
@@
Figure 4: Illustrations of Combination and Transformation function in Transformer architecture models.
@@ 4.2.1 Mutual Information
We adopted the mutual information estimate method proposed by  (Voita et al., 2019), which mainly involves converting the representations into discrete variables through clustering, and then calculating the mutual information. For specific details, please refer to Appendix E.
@@
Figure 5: Contribution ratio analysis based on Mutual Information (MI) for the vanilla and FFN-Wider models.
@@ 4.2.2 Token Prediction
Figure 6: Contribution ratio analysis based on Token Prediction (TP) for the vanilla and FFN-Wider models.
@@ 5 Combination Adjustable Architecture
Although we observed the FFN layer (a transformation function) in FFN-Wider Transformers has a higher contribution ratio, we cannot assert that this is the reason for the degradation in the model’s base capabilities. This may merely be a correlation rather than a decisive factor in base capabilities.
@@ 5.1.2 Direct Pathway in MHA
Figure 8: The trends of contribution ratio and base capability as the width ratio of the Outer-FFN changes.
@@ 5.2.1 Trend of Contribution Ratio
We conducted experiments on large-scale (H=768) BERT and GPT models. Initially, a series of models with different width ratios were pre-trained, ensuring their pre-training performance was aligned. Then, we calculated the contribution ratio of Outer-FFN using Token Prediction (TP) and plotted these trends, as shown in Figures 8(a) and 8(c).
@@ 5.2.2 Trend of Base Capability
We tested the performance of each model above in out-of-distribution (OOD) language modeling and plotted the trend in Figures 8(b) and 8(d).
@@
However, this overall trend does not mean the contribution of transformation function should be reduced to zero. A minor contribution from transformation function might still be crucial, as indicated by the sole anomalous result in GPT. BERT was unaffected, which is related to bidirectional attention still being able to leak few transformation contribution. More details are in Appendix F.
Table 1: Experimental results of various BERT models.
Model Pile GLUE SGLUE Others
(22 parts) (8 tasks) (8 tasks) (6 tasks)
(H=128,  Pre-Training $\bm{\approx}$ –2.61)
Vanilla BERT -3.211 78.45
FFN-Wider BERT -3.256 76.66
FFN-Wider BERT w/ CEA -3.202 78.20
   - w/o Direct Pathway in MHA -3.230 77.29
(H=768,  Pre-Training $\bm{\approx}$ –1.53)
Vanilla BERT -2.158 83.89 64.41 52.14
FFN-Wider BERT -2.196 82.60 62.82 51.61
FFN-Wider BERT w/ CEA -2.149 83.86 64.22 53.12
Table 2: Experimental results of various GPT models.
Model Pile 0-Shot 1-Shot
   (22 parts)    (9 tasks)    (9 tasks)
(H=128,  Pre-Training $\bm{\approx}$ –4.06)
Vanilla GPT -4.706
FFN-Wider GPT -4.728
FFN-Wider GPT w/ CEA -4.685
(H=768,  Pre-Training $\bm{\approx}$ –3.19)
Vanilla GPT -3.878 45.82 46.09
FFN-Wider GPT -3.911 44.60 44.84
FFN-Wider GPT w/ CEA -3.882 45.52 45.84
Table 3: The MoE results. AX: ArXiv, DM: DM Mathematics, GH: Github, SE: Stack Exchange, UI: Ubuntu IRC.
Model AX DM GH SE UI MRPC STS-B
(H=128,  Pre-Training $\bm{\approx}$ –2.64)
Vanilla BERT   -3.324   -2.806   -3.343   -3.407   -4.005   83.06   88.15   83.35   83.33
MoE BERT -3.445 -2.930 -3.450 -3.473 -4.074 79.78 85.52 81.87 81.53
MoE BERT w/ CEA -3.392 -2.842 -3.337 -3.402 -4.086 79.70 85.84 83.53 83.26
(H=768,  Pre-Training $\bm{\approx}$ –1.51)
Vanilla BERT -2.275 -2.184 -2.351 -2.369 -3.001 87.35 90.74 88.12 87.95
MoE BERT -2.341 -2.261 -2.420 -2.410 -3.075 84.15 89.08 85.65 85.31
MoE BERT w/ CEA -2.271 -2.211 -2.399 -2.417 -3.040 85.29 89.76 85.55 85.20
@@ 6.1 Width Ratio Selection
This work is interested in how changes in architecture can maximally enhance base capabilities. Therefore, we selected the optimal width ratio for Combination Enhanced Architecture (CEA).
The Outer-FFN width ratio for FFN-Wider BERT w/ CEA is set to 0%, and for FFN-Wider GPT w/ CEA, it is set to 12.5%.
@@ 6.2 Further Experiments
For BERT, we tested fine-tuning performance on GLUE, SuperGLUE and other datasets. For GPT, we tested zero-shot and one-shot performance on multiple datasets. For both, we tested OOD language modeling on Pile. For small-scale models, we removed some experiments beyond their capabilities. The brief results are shown in Table 1 and 2, and the detailed results are in Appendix G.
@@
Additionally, we conducted similar experiments for other width ratios that can align pre-training performance and steps simultaneously, which can indisputably prove improvements come from architecture, as detailed in Appendix H.
7 From FFN-Wider Transformers to MoE Architecture Transformers
In this section, we extend our explanation and CEA to Mixture of Experts (MoE) architecture Transformers Lepikhin et al. (2021); Fedus et al. (2022).

### v3

# How does Architecture Influence the Base Capabilities of Pre-trained Language Models? A Case Study Based on FFN-Wider and MoE Transformers

@@ Abstract
Pre-trained language models have been proven to possess strong base capabilities, which not only excel in in-distribution language modeling but also show powerful abilities in out-of-distribution language modeling, transfer learning and few-shot learning. Unlike existing work focusing on the influence of scale on base capabilities, our work examines the influence of architecture on those. Specifically, our concern is: How does architecture influence the base capabilities of pre-trained language models? In this work, we attempt to explain and reverse the decline in base capabilities caused by the architecture of FFN-Wider Transformers, seeking to provide some insights. Through analysis, we found the contribution ratio of Multi-Head Attention (a combination function) to pre-trained language modeling is a key factor affecting base capabilities. FFN-Wider Transformers reduce the contribution ratio of this combination function, leading to a decline in base capabilities. We confirmed this by experiments and proposed Combination Enhanced Architecture (CEA) to address the decline in base capabilities of such models. Significantly, we extended our explanation and CEA to Mixture of Experts (MoE) Transformers. We successfully achieved significant improvements in base capabilities on a 14B parameter MoE model, demonstrating the practical application value of our work. This also indicates that our analysis has a certain guiding significance for architecture analysis, architecture improvement and architecture design.
@@ 1 Introduction
Recent research has discovered that pre-trained language models possess strong base capabilities [24, 7, 3, 22]. They can not only address in-distribution language modeling which is usually their pre-training objective, but also unexpectedly excel in out-of-distribution language modeling, transfer learning, few-shot learning, etc. This has attracted the attention of many researchers.
@@
Figure 1: Illustration showing that: 1) the synchronous improvement in model base capability as the contribution ratio of the Outer-FFN layer (a transformation function) decreases, that is, the contribution ratio of the MHA layer (a combination function) increases. This reveals a key factor affecting model’s base capabilities. 2) Combination Enhanced Architecture (CEA) was designed based on this factor and applied to MoE models, resulting in an improvement in base capability.
In this process, much attention has been focused on analyzing scale, leading to the formulation of compelling scaling laws [14, 11] that drive the trend of enhancing base capability by increasing parameter numbers, data volume and training tokens. In contrast, the impact of architecture has not received sufficient attention. According to the basic principle of inductive bias in machine learning, model architecture is also a crucial factor affecting base capabilities, and its impact could be equally decisive.
Some studies have already noted the significant influence of architecture on the base capabilities. For instance,  [31] have observed considerable differences in base capabilities among various Transformer architecture variants. Some variants, though larger in scale and more powerful in pre-training performance than vanilla Transformers, exhibit significantly reduced performance in downstream tasks. This suggests different architecture models vary greatly in converting pre-training performance into base capabilities. Simply increasing scale does not resolve all issues, and exploring the impact of architecture on base capabilities is crucial.
@@
Specifically, we first focus on FFN-Wider Transformers. This architecture has wider FFN layers, but we found that the change leads to a significant decrease in base capabilities compared to vanilla Transformers. As shown in Figure 2, under similar pre-training performance, the FFN-Wider models exhibit a noticeable decline in base capabilities compared to the vanilla models. We believe such a simple change, leading to significant differences in base capabilities, makes for a good object of study in exploring how architecture impacts base capabilities.
@@
Then, we attempted to explain this change in base capabilities. Through analysis, we concluded that the MHA (Multi-Head Attention) layer in the Transformer is a combination function, and the FFN (Feed-Forward Network) layer is a transformation function, with the former being a focused expression of the combinability of language. During their contribution to pre-trained language modeling, the actual contribution ratio of the MHA layer, as a combination function, is a key factor affecting the model’s base capabilities. The FFN-Wider Transformer models directly enhance the FFN layer, indirectly reducing the combination function’s actual contribution ratio to pre-trained language modeling, thereby leading to a significant decline in base capabilities.
@@
To validate our explanation for it, we designed a Combination Adjustable Architecture (CAA), as depicted in Figure 1(a). This architecture bifurcates a wider FFN into two parts with adjustable width ratios: one remains in its original position as a transformation function, known as the Outer-FFN, and the other is relocated within the MHA layer, transformed through a special design into an Inner-FFN that solely enhances the combination function. We controlled the width ratio of the Outer-FFN, reducing it gradually from 100% to 0%. In Figure 1(b), we observed that the actual contribution ratio of the Outer-FFN progressively decreased, indicating a corresponding increase in the actual contribution ratio of the MHA layer. At the same time, we also observed a gradual improvement in base capabilities. These reveal a key phenomenon that confirms our explanation: as the actual contribution ratio of the MHA layer (a combination function) increases, there is a general synchronous improvement in the model’s base capabilities.
@@
Subsequently, we identified the optimal width ratio of two parts of FFN for the Combination Adjustable Architecture (CAA) and established it as Combination Enhanced Architecture (CEA), and comparing this architecture with the original FFN-Wider Transformer. We conducted experiments on various scales of BERT and GPT models, with all results robustly supporting our explanation.
Importantly, we also noticed that existing work has observed a similar decline in base capabilities in the Mixture of Experts (MoE) Transformers. We applied our explanation and CEA to MoE Transformers as well. We pre-trained a 14B parameter GPT architecture MoE model and its improved version with CEA on 100B tokens (using CC and C4 from SlimPajama [29]). With the same number of parameters, computational load, and pre-training steps for both versions, the improved version with CEA showed significant improvements in out-of-distribution language modeling and few-shot learning. The performance of the out-of-distribution language modeling (average performance across all domains in SlimPajama except for CC and C4) is shown in Figure 1(c), which can prove the practical application value of our work.
Overall, the actual contribution ratio of the MHA layer (a combination function) is likely a universal factor affecting model’s base capabilities, which can provide valuable insights for architecture analysis, improvement and design.
@@ 2 Background
Subsequent sections will involve specific analyses and experiments of FFN-Wider and MoE transformers. Therefore, in this section, we introduce the relevant background of base capabilities, as well as the unified evaluation schemes and evaluation tasks.
@@ 2.1 Base Capabilities
Transfer Learning   This is a recognized base capability of pre-trained language models. The works proposing GPT [24] and BERT [7] have established the "pre-training and fine-tuning" transfer learning paradigm under the Transformer architecture.
Few-shot Learning   This is also a recognized base capability. Some works [25, 3] found large-scale pre-trained language models could complete many NLP tasks with no or only a few demonstrations, revealing their remarkable few-shot learning capabilities.
@@ 2.2 Evaluation Schemes
The experiment with MoE transformers is an exception, because we need to demonstrate practicality. Therefore, in the comparative experiments between the vanilla MoE model and the improved MoE model, not only is pre-training performance aligned, but pre-training steps, parameter numbers, and computational load are also kept consistent.
2.3 Evaluation Tasks
Due to subsequent experiments involving out-of-distribution language modeling evaluation, transfer learning evaluation and few-shot learning evaluation, it is necessary to clarify the selection of pre-training corpus and out-of-distribution test corpus, as well as the selection of downstream tasks. We provide a detailed introduction in Appendix B.
@@ 2.3 Evaluation Tasks
We found that, at the same level of pre-training performance, the Transformer models with wider FFN layers exhibit a noticeable decline in performance on most downstream tasks, indicating a deterioration in their base capabilities. This presents us with a good research object.
Figure 2: Comparison of the base capabilities between FFN-Wider and Vanilla Transformers.
@@ 4.1 Combination and Transformation
Transformers consist of MHA and FFN layers. Considering a certain position in a sequence, the updated representation obtained after an MHA layer is a combination of the entire sequence context; whereas the updated representation obtained after an FFN layer is from a transformation of that position representation alone, context-insensitive.
@@ 4.2.1 Mutual Information
We adopted the mutual information estimate method proposed by  [32], which mainly involves converting the representations into discrete variables through clustering, and then calculating the mutual information. For the definition of MI and more specific details, please refer to Appendix F.
@@ 4.2.2 Token Prediction
[…]
Figure 3: Contribution ratio analysis based on Mutual Information(MI) for various transformers.
[…]
Figure 4: Contribution ratio analysis based on Token Prediction (TP) for various transformers.
@@ 5 Combination Adjustable Architecture
Although the FFN layer in FFN-Wider Transformers has a higher contribution ratio, we cannot assert this is the reason. This may merely be a correlation rather than a decisive factor in base capabilities.
@@ 5.1.2 Direct Pathway in MHA
[…]
Figure 5: Overview of our proposed Combination Adjustable Architecture (CAA).
[…]
Figure 6: Outer-FFN contribution ratio and base capability under different width ratios.
@@ 5.2.1 Trend of Contribution Ratio
We conducted experiments on large-scale (H=768) BERT and GPT models. Initially, a series of models with different width ratios were pre-trained, ensuring their pre-training performance was aligned. Then, we calculated the contribution ratio of Outer-FFN using Token Prediction (TP) and plotted these trends, as shown in Figures 6(a) and 6(c).
@@ 5.2.2 Trend of Base Capability
We tested the performance of each model above in out-of-distribution (OOD) language modeling and plotted the trend in Figures 6(b) and 6(d).
@@
However, this overall trend does not mean the contribution of transformation function should be reduced to zero. A minor contribution from transformation function might still be crucial, as indicated by the sole anomalous result in GPT. BERT was unaffected, which is related to bidirectional attention still being able to leak few transformation contribution. More details are in Appendix G.
[…]
Model Pile GLUE SGLUE Others (22 parts) (8 tasks) (8 tasks) (6 tasks) (H=128,  Pre-Training $\bm{\approx}$ –2.61) Vanilla BERT -3.211 78.45 FFN-Wider BERT -3.256 76.66 FFN-Wider BERT w/ CEA -3.202 78.20      - w/o Direct Pathway in MHA -3.230 77.29 (H=768,  Pre-Training $\bm{\approx}$ –1.53) Vanilla BERT -2.158 83.89 64.41 52.14 FFN-Wider BERT -2.196 82.60 62.82 51.61 FFN-Wider BERT w/ CEA -2.149 83.86 64.22 53.12
Figure 7: The results of various BERT models.
[…]
Model Pile 0-Shot 1-Shot    (22 parts)    (9 tasks)    (9 tasks) (H=128,  Pre-Training $\bm{\approx}$ –4.06) Vanilla GPT -4.706 FFN-Wider GPT -4.728 FFN-Wider GPT w/ CEA -4.685 (H=768,  Pre-Training $\bm{\approx}$ –3.19) Vanilla GPT -3.878 45.82 46.09 FFN-Wider GPT -3.911 44.60 44.84 FFN-Wider GPT w/ CEA -3.882 45.52 45.84
Figure 8: The results of various GPT models.
@@ 6.1 Width Ratio Selection
We selected the optimal width ratio for Combination Enhanced Architecture (CEA). The Outer-FFN width ratio for FFN-Wider BERT w/ CEA is set to 0%. For FFN-Wider GPT w/ CEA, it is 12.5%.
@@ 6.2 Further Experiments
For BERT, we tested fine-tuning performance on GLUE, SuperGLUE and other datasets. For GPT, we tested zero-shot and one-shot performance on multiple datasets. For both, we tested OOD language modeling on Pile. For small-scale models, we removed some experiments beyond their capabilities. The brief results are shown in Table 8 and 8, and the detailed results are in Appendix H.
@@
Additionally, we conducted similar experiments for other width ratios that can align pre-training performance and steps simultaneously, which can indisputably prove improvements come from architecture, as detailed in Appendix I.
Metric # Shot Vanilla GPT 1.3B Vanilla MoE 14B MoE 14B w/ CEA
# Total Params N/A 1.3B 14B 14B
# Activated Params N/A 1.3B 1.5B 1.5B
# Total Experts N/A - 64 64
# Activated Experts N/A - 1 1
# Training Tokens N/A 100B 100B 100B
SlimPajama-CC&C4 (Loss) N/A 2.382 2.315 2.303
SlimPajama-Arxiv (Loss) N/A 2.353 2.320 2.239
SlimPajama-Book (Loss) N/A 2.796 2.761 2.585
SlimPajama-Github (Loss) N/A 1.989 1.989 1.840
SlimPajama-Stack (Loss) N/A 2.653 2.588 2.465
SlimPajama-Wiki (Loss) N/A 3.137 3.155 2.964
LAMBADA (PPL) 0-shot 23.7 24.8 19.3
LAMBADA (Acc.) 0-shot 36.6 36.4 39.7
MMLU (Acc.) 5-shot 30.8 30.6 31.4
OpenBookQA (Acc.) 5-shot 36.6 35.7 37.2
ARC Easy (Acc.) 5-shot 57.5 56.8 59.2
ARC Challenge (Acc.) 5-shot 31.6 30.0 32.3
BoolQ (Acc.) 5-shot 62.2 62.2 62.1
RACE Middle (Acc.) 5-shot 42.3 44.4 45.5
RACE High (Acc.) 5-shot 36.1 36.0 36.6
SIQA (Acc.) 5-shot 41.0 41.7 43.7
SCIQ (Acc.) 5-shot 78.1 72.7 84.1
HellaSwag (Acc.) 5-shot 43.7 45.8 47.5
COPA (Acc.) 5-shot 69.2 69.2 70.4
PIQA (Acc.) 5-shot 70.4 70.6 71.5
StoryCloze (Acc.) 5-shot 67.7 68.0 69.5
WinoGrande (Acc.) 5-shot 52.7 53.1 53.4
Winograd (Acc.) 5-shot 67.2 66.7 66.8
Table 1: The results of Vanilla GPT 1.3B, Vanilla MoE 14B and MoE 14B w/ CEA. indicates the test set drawn from the same distribution as the pre-training data.
7 From FFN-Wider Transformers to MoE Transformers